You are here: Foswiki>Commissioning Web>Com20100630 (30 Jun 2010, PascaleHibon)Edit Attach

Commissioning: 30 June 2010 UT

Observer: Pascale H, Kate R, Dave T (LBTO)
Telescope Operator: Steve A (LBTO)
Telescope Support: (LBTO)
Instrument Support: Tim S (LBTO)
Adaptive Secondary Operator: Kate R, Pascale H (LBTO), Marco X (Firenze)

Summary

Short night due to weather.
Pointings mainly for Astigmatism test for Marco and to test also the Pointing.
Problem with shutter doors. See Issue Track 2498.

Preparations

03:35 Rain clouds around the telescope. Risk of rain.
The Shell of the M2 has been set and is on Flat mode with Offset OFF

Open

06:27 open Humidity=72% T_air=10C Wind=5-6m/s

Details

06:36 pointto M5_0555 Tel="TRACK" gs=0 PA=0
06:41 pointto M5_0555 Tel="ACQUIRE" gs=0 PA=0 PAM-
CA=7 IE=-37 06:46 pointto BS9169 Tel="ACTIVE" gs=0 Elevation 66deg
Seeing~2"

Test for Marco + Pointing

  • The test for Marco consists on pointing on different stars at different elevation (between 30 and 40 deg) with a reset of the Active Optices Corrections between each pointing. This test will allow us to see if the presence of astigmatism is correlated with gravity and if it appears only at a certain elevation.

  • We also took advantage of this test to test the pointing of the telescope by going from low to high elevation.

06:58 pointto BS9153 Tel="ACTIVE" gs=0 Elevation 30deg
07:01 pointto BS9153 Tel= "ACQUIRE" gs=0 PA=0 PAM- -> Strong Astigmatism
Steve is running a ptreset and a ptspiral.
CA=-7 IE=-51
07:14 pointto ACT0369 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 70deg
CA= 5 IE= -41
07:22 pointto ACT0317 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 88deg -> A little of Astigmatism
CA=8 IE=-35
07:26 pointto ACT0096 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 25deg -> Astigmatism of 6arsec long
IA= -413
07:33 pointto ACT0317 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 89deg -> Astigmatism of ~1arcsec long
CA= 12 IE=-43
07:50 Reset Active Optics Correction
CA= 6 IE= -46
07:53 pointto ACT0317 Tel="ACTIVE" gs=0 Elevation 85deg
No Collimation possible. This source is saturated on the WFS
07:59 pointto BS9160 Tel="ACTIVE" gs=0 Elevation 30deg
Star not found
08:02 pointto BS9160 Tel="ACQUIRE" gs=0 PA=0 PAM-
Steve is doing a ptspiral
CA=15 IE=-75
08:06 pointto BS9160 Tel="ACTIVE" gs=0 Elevation 30deg -> Astigmatism of 7arcsec
08:07 pointto BS9188 Tel="ACTIVE" gs=0 Elevation 40deg

Close

08:11 Close the telescope Humidity=92% Dewpoint=8C
10:55 the shutter doors are finally closed. Humidity=98% Dewpoint 8C (Issue Track 2498)
Topic revision: r3 - 30 Jun 2010, PascaleHibon
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback